# Integrability in Perturbed Black Holes: Background Hidden Structures

### v1

@@ I Introduction: Hidden symmetries in black hole dynamics
A specially important instance of a structurally stable dynamical feature in our context is given by the so-called soliton resolution conjecture [4], according to which generic global-in-time (general relativistic) non-linear wave dynamics would decouple universally at late times into “soliton” solutions plus “radiation”. A dramatic example of such soliton resolution behaviour in GR would be provided by the so-called “establishment picture of gravitational collapse”, namely a sequence of theorems and conjectures that provide a universal and structural stable picture in which generic gravitationally collapsing configurations asymptotically approach a universal class of BH spacetimes (namely sub-critical Kerr, the soliton solution in this setting) plus the (gravitational) radiation emitted at null infinity and through the horizon.
More broadly, we aim at exploring the role of such a class of structurally stable features of the theory in the setting of general BH dynamics, including binary BH (BBH) mergers. The soliton resolution conjecture properly applies to non-linear dispersive equations [5]. In this setting, as a methodological approach, we adopt the perspective recently proposed in [6] (see also [7, 8]). Such a methodology relies on two elements, on the one hand a “wave-mean flow” approach to the dynamics   [^1] More specifically, this corresponds to the layer in section 5 of the hierarchical proposal to BBH dynamics in Ref. [6] (see Table 1 in that reference and also Refs. [9, 7]). and, on the other hand, a so-called “asymptotic reasoning” strategy [10]. The former proposes the effective separation of the dynamics into “fast” and “slow” degrees of freedom (DoFs), whereas the latter filters away some DoFs of the theory thus sacrificing precision and exactness in order to unveil the underlying (universal) structurally stable patterns, typically in an asymptotic limit of a relevant physical parameter.
@@ I.1 Wave-mean flow and asymptotic reasoning
    Linear scattering of “test fields” on dynamical background: “fast” DoFs do not backreact on the dynamically slow background.
@@
The systematic study of such a hierarchy defines a work program, which by itself constitutes a subprogram of the one described in Refs. [6, 7], each layer focusing on a particular dynamical aspect of the problem. The main goal of this work is to focus on the identification of the relevant algebraic structures controlling the structurally stable aspects of the dynamics, more specifically, on the ones of the bulk spacetime (slow DoFs). Pushing a bit further the asymptotic reasoning scheme, the Ansatz is that such algebraic structures, meant to be present and controlled by the whole hierarchy, are more easily identifiable in the lowest layer —point iii) above. In this article, we focus on the linear scattering on a stationary background (the BH background) as a first step in a further reaching program.
@@ I.2 Background structures in linear scattering
Moreover, the structure of the space of master functions and equations gets enriched by the introduction [14] of inverse scattering techniques [21, 22, 23, 24, 25]. These techniques can be used to solve certain non-linear evolution problems [26, 27] by establishing a mathematical connection with linear wave scattering problems. Those non-linear PDEs that can be solve by these methods are then characterised as integrable. The Korteweg-de Vries (KdV) equation, which describes a very wide range of physical phenomena [28], provides a paradigmatic example. The key idea behind the application of the inverse scattering method to this integrable PDE setting is to cast the unknown satisfying the non-linear evolution PDE as a time-evolving potential that can be associated to a (linear) scattering problem. That is, the unknown can then be reconstructed from the (non-linear) evolution of the associated scattering data. The scattering data is obtained from the associated linear, and time-independent, problem. In the case of the KdV equation, the linear problem is the time-independent Schrödinger equation. Another remarkable result is the appearance of an infinite series of conservation laws for the KdV equation with the corresponding set of conserved quantities [29, 30, 31], the so-called KdV integrals. We can apply these techniques to our master equations in the frequency domain and generate the associated set of KdV integrals [14]. One can see that the KdV equation constitutes an isospectral deformation of the master equations, in such a way that the transmission coefficients and the QNMs are preserved by the KdV deformation. Crucially, it was shown in [14] that the KdV integrals are invariant under Darboux transformations. That is, they are the same for all the potentials in the set of master equations.
@@ I.3 Hyperboloidal approach to BHPT
The use of hyperboloidal foliations have proved very successful in its application to both fundamental and applied problems in spacetime dynamics and BH scattering. It provides one of the main ingredients in key results such as the semi-global stability of Minkowski by Friedrich [35], BH stability results in cosmological scenarios [36, 37, 38], the characterisation of QNMs as eigenvalues of a non-selfadjoint operator [39, 40], the numerical stability of the exploration of solution resolution in [4], or the analysis of the spectral instability of BH QNMs [41, 42]. In recent years hyperbolodial foliations have become one of the main tools for the theoretical and numerical analysis of compact object sources of gravitational waves (cf. excellent discussions in [43, 44, 45, 46, 47]).
A most important aspect of the hyperboloidal scheme in our present setting, namely focusing on the identification of the relevant algebraic structures in BH scattering, is that it provides a neat separation of bulk and asymptotic degrees of freedom. This feature manifests explicitly in the structure of the infinitesimal time generator of perturbations and it is the consequence of two ingredients: i) the choice of a “hyperboloidal time” $\tau$ related to the “Cauchy time” $t$ through an appropriate height function (i.e. $\tau=t+h(x)$) introducing the hyperboloidal foliation $\{\Sigma_{\tau}\}$ to cast the dynamics, and ii) a compactification of the hyperboloidal hypersurfaces $\Sigma_{\tau}$. The first element adapts the treatment to the relevant outgoing asymptotics, whereas the second one recasts such outgoing boundary conditions in terms of the regularity of the scattered field at the boundaries (the latter is akin  [^2] We thank J.A. Valiente-Kroon for insisting on this point. We refer to [48] for a discussion on the adaptation of Melrose’s “geometric scattering” to the hyperboloidal approach to spacetime dynamics. to the underlying philosophy in Melrose’s ‘geometric scattering’ [49]). Enforcing points i) and ii) casts the wave dynamics of the master function $\phi$ in Schrödinger form
@@
The structure of $\mathbb{L}$ reflects the above-commented separation into bulk and asymptotic structures, respectively encoded in the ${\cal L}_{1}$ and ${\cal L}_{2}$ parts. On the one hand, the operator ${\cal L}_{1}$ stands as the structural counterpart in the hyperboloidal scheme of the Schrödinger operator in the Cauchy slicing setting of section I.2. Specifically, ${\cal L}_{1}$ is a (singular) Sturm-Liouville operator, namely a second-order elliptic operator that includes the effective scattering potential and provides the elements for the discussion of the background structures in linear scattering (namely, Darboux covariance and KdV structures). Its “singular” character, namely the vanishing at the boundaries of its principal as a consequence of the (‘Melrose’s) compactification of $\Sigma_{\tau}$, results in the absence of explicit boundary conditions that are now encoded in the structure of the operator $\mathbb{L}$ itself. On the other hand, the operator ${\cal L}_{2}$ is the structural piece encoding the outgoing nature of the boundary conditions of the scattering problem. That is, if ${\cal L}_{1}$ incorporates the boundary conditions into $\mathbb{L}$, it is ${\cal L}_{2}$ the element that characterises precisely its outgoing nature. It can actually be shown [41] that ${\cal L}_{2}$ is the term responsible of the non-selfadjoint character of $\mathbb{L}$ and that the corresponding dissipative character of the latter is precisely due to the flow of the scattered field through the spacetime boundaries, that is controlled by the parameters in ${\cal L}_{2}$. Moreover, the term breaking the selfadjoint nature of the infinitesimal time generator, i.e. the operator $\mathbb{L}-\mathbb{L}^{\dagger}$ (where the adjoint $\mathbb{L}^{\dagger}$ is defined with respect to a natural ‘energy scalar product’) has a distributional nature with support at null infinity and the BH horizon. In other words, the missing DoFs in the hyperboloidal non-selfadjoint formulation of the problem, associated with the loss of DoFs propagating from the bulk through the boundary, are sharply accounted by the flux controlled by the ${\cal L}_{2}$ term, the latter being indeed a structure essentially living on the boundaries. This has led to the proposal in [50] that one can account for the loss of “fast” DoFs reaching the boundaries in terms of formal DoFs “living” on the spacetime boundaries, that would work as a kind of “geiger counters” of “fast” DoFs (their inclusion in the problem rendering it conservative, since the total account of energy is conserved by construction). More ambitiously, Ref. [50] proposes that asymptotic symmetries [namely, the Bondi-Metzner-Sachs (BMS) group of symmetries [51, 52], or an appropriate extension] would stand as a dynamical symmetry generating the phase space of DoF on the boundary, namely the (asymptotic) “fast” DoFs.
@@ I.4 The “KdV-Virasoro-Schwarzian derivative” triangle: bulk structures in a BH dynamics program
Dwelling then in the bulk part, this work focuses on the bulk symmetries and explores the heuristic intuition that the KdV structure of BHPT [14, 15] is related to conformal symmetries of the bulk DoFs. This leads to identification of the role of the Virasoro algebra in this setting and that of the Schwarzian derivative in the transformation of the scattering potential under conformal mappings. In this setting, the “KdV-Virasoro-Schwarzian derivative” triangle is identified as providing the basic ingredients for the underlying bulk integrability (cf. [53] for a systematic discussion in a larger integrability setting). This is done first in the setting of a Cauchy description, where the KdV symmetries first appeared [14, 15], then finding that such a “bulk” symmetry structure is (partially) preserved when passing to the hyperboloidal foliation.
The plan of the article is the following. In Sec. II we first review the basic concept and tools of BHPT and then introduce the complete landscape of the possible master equations decouplings allowed by Einstein equations. In particular, in subsection II.3 we review the symmetry structure hiding behind such infinite hierarchy of master equations, namely Darboux covariance. In Sec. III we explore the bulk symmetries, in Cauchy slices, associated to KdV integrable structures. First, we introduce the KdV isospectrality and some interesting relations to SUSY quantum mechanics. Then, we show the connection with the conformal transformations and the Virasoro algebra. We conclude this section showing how the conformal transformations acts on the master equations. In Sec. IV we present how the structures studied in the previous section can be carried to a hyperboloidal scheme. For this we review the hyperboloidal approach in BHPT and introduce a general transformation of the master function, showing the covariance of the hyperboloidal master equation under such transformation. Finally, we conclude with a specific choice of the transformation which provides the connection with the bulk symmetries investigated in Sec. III.
@@ II Linear scattering on a stationary background: BHPT
where one assumes that $|h_{\mu\nu}|\ll|\widehat{g}_{\mu\nu}|$ (see, e.g. [57, 58] for a precise formulation of this statement). Here, the background metric $\widehat{g}_{\mu\nu}$ satisfies the full Einstein equations while the dynamic of the perturbations $h_{\mu\nu}$ is linear. Schematically, this can be written as follows
@@
where $\delta G$ is the linearized Einstein operator. However, the identification background/“slow” and perturbative/“fast” DoFs is not apparent at this stage, as we still have to identify the correct background quantities which are related to integrable/stable structures and the perturbative DoFs associated to the linear asymptotic dynamics. The perturbation theory framework guides us towards the decoupling of Einstein equations for the perturbations (7) into (infinite) master equations with a potential and a wave function which will be directly related to “slow” and “fast” DoFs respectively. This already reduces the problem to the study of linear scattering (of gravitational perturbations) on a stationary background  [^3] We however only consider static spherically symmetric backgrounds in this work..
@@ II.1 The frozen slow DoFs: spherically symmetric background
where $f(r)$ is given in Eq. (9) and we have chosen $x^{a}=(t,r)$ and $\Theta^{A}=(\theta,\varphi)$. In Eq. (10), $r(x^{a})$ is the area radial coordinate and $r^{2}$ is the geometry warp factor. The covariant derivative on $S^{2}$ is denoted by a vertical bar ($\Omega_{AB|C}=0$) while for $M^{2}$ we use a colon ($g_{ab:c}=0$). The volume Levi-Civita tensor in $S^{2}$ is denoted by $\epsilon_{AB}$ while in $M^{2}$ is denoted by $\varepsilon_{ab}\,$. Something interesting about the metric of $M^{2}$ [see Eq. (11)] is that, if we use the tortoise coordinate
@@
It is important to stress that while the perturbative approach linearizes the dynamics of fast degrees of freedom, it provides a simplified but non-trivial framework to investigate the presence of (universal) integrable structure governing the slow/background DoFs. Indeed, while the dynamic of perturbations is linear by construction [see Eq. (7)], the background is a fully nonlinear solution with a rich structure acting on the perturbations.
@@ II.2 Linear dynamics of the fast DoFs: master equations
We are not going to show explicit expressions for the harmonic components of the metric perturbations since they are not relevant in this work. They can be found, for instance, in Refs. [14, 55]. We also drop the harmonic indices from now on for the sake of clarity. The separation in Eq. (16) allows to decouple Einstein perturbative equations into an infinite set of equations for the different $(\ell,m)$ and parity modes (see, e.g. [61, 62]). Now, let us consider the Einstein perturbative equations for just a single mode, that is, for fixed harmonic numbers $(\ell,m)$ and parity. These equations, seven for the even-parity case and three for the odd-parity case, are $1+1$ linear PDEs (i.e. only with dependence on the coordinates of $M^{2}$) that can be combined in such a way that we can isolate (decouple) special gauge-invariant combinations of the metric perturbations (including their first-order derivatives), known as master functions. In this way, the perturbative Einstein equations for a single mode reduce to wave-type equations with a potential, known as master equations. The concrete form of the master equations is
@@
where $x$ is the tortoise coordinate introduced in Eq. (13). In the wave equation (17), $\Psi^{\rm even/odd}(t,r)$ is the even/odd-parity master function; and $V^{\rm even/odd}_{\ell}(r)$ is the associated potential, which only depends on the harmonic number $\ell$ and the parity. For the cases where we have explicit expressions of the potentials, these are in terms of the radial area coordinate $r$, but for the purposes of solving the master wave-type equations, we have to understand this dependence on $r$ as on a function of the radial coordinate, i.e. on $r=r(x)$.
@@
where $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are arbitrary constants. Therefore, for odd-parity perturbations, the master function is a linear combination of the Regge-Wheeler (RW) [18] and the Cunningham-Price-Moncrief (CPM) [63, 64, 65] master functions. In the case of even-parity perturbations, the most general master function is instead a linear combination of the Zerilli-Moncrief (ZM) [19, 66] and another one which was first found in covariant form in Ref. [13]  [^5] This master function already appeared in Ref. [82], in Schwarzschild coordinates and in the Regge-Wheeler gauge, as a better choice for computations in the context of the close limit approximation.. We are not going to show here the explicit covariant expressions as they can be found in Refs. [13, 17].

### v2

@@ I Introduction: Hidden symmetries in black hole dynamics
A specially important instance of a structurally stable dynamical feature in our context is given by the so-called soliton resolution conjecture [4], according to which generic global-in-time (general relativistic) non-linear wave dynamics would decouple universally at late times into “soliton” solutions plus “radiation”. A dramatic example of such soliton resolution behaviour in GR would be provided by the so-called “establishment picture of gravitational collapse”, namely a sequence of theorems and conjectures that provide a universal and structural stable picture in which generic gravitationally collapsing configurations asymptotically approach a universal class of BH spacetimes (namely sub-critical Kerr, the soliton-like solution in this setting) plus the (gravitational) radiation emitted at null infinity and through the horizon.
More broadly, we aim at exploring the role of such a class of structurally stable features of the theory in the setting of general BH dynamics, including binary BH (BBH) mergers. The soliton resolution conjecture properly applies to non-linear dispersive equations [5]. In this setting, as a methodological approach, we adopt the perspective recently proposed in [6] (see also [7, 8, 9]). Such a methodology relies on two elements, on the one hand a “wave-mean flow” approach to the dynamics   [^1] More specifically, this corresponds to the layer discussed in Sec. V of the hierarchical proposal to BBH dynamics in Ref. [6] (see Table 1 in that reference and also Refs. [10, 7, 9]). and, on the other hand, a so-called “asymptotic reasoning” strategy [11]. The former proposes the effective separation of the dynamics into “fast” and “slow” degrees of freedom (DoFs), whereas the latter filters away some DoFs of the theory thus sacrificing precision and exactness in order to unveil the underlying (universal) structurally stable patterns, typically in an asymptotic limit of a relevant physical parameter.
@@ I.1 Wave-mean flow and asymptotic reasoning
    Linear scattering of “test fields” on a dynamical background: “fast” DoFs do not backreact on the dynamically slow background.
@@
The systematic study of such a hierarchy defines a work program, which by itself constitutes a subprogram of the one described in Refs. [6, 7] (see also [9]), each layer focusing on a particular dynamical aspect of the problem. The main goal of this work is to focus on the identification of the relevant algebraic structures controlling the structurally stable aspects of the dynamics, more concretely, on the specific ones of the bulk background spacetime (slow DoFs). Pushing a bit further the asymptotic reasoning scheme, the Ansatz is that such algebraic structures, meant to be present and controlling the whole hierarchy, are more easily identifiable in the lowest layer —point iii) above. In this article, we focus on the linear scattering on a stationary background (the BH background) as a first step in a further reaching program.
@@ I.2 Background structures in linear scattering
Moreover, the structure of the space of master functions and equations gets enriched by the introduction [15] of inverse scattering techniques [22, 23, 24, 25, 26]. These techniques can be used to solve certain non-linear evolution problems [27, 28] by establishing a mathematical connection with linear wave scattering problems. Those non-linear PDEs that can be solved by these methods are then characterised as integrable. The Korteweg-de Vries (KdV) equation, which describes a very wide range of physical phenomena [29], provides a paradigmatic example. The key idea behind the application of the inverse scattering method to this integrable PDE setting is to cast the unknown satisfying the non-linear evolution PDE as a time-evolving potential that can be associated to a (linear) scattering problem. That is, the unknown can then be reconstructed from the (non-linear) evolution of the associated scattering data. The scattering data is obtained from the associated linear, and time-independent, problem. In the case of the KdV equation, the linear problem is the time-independent Schrödinger equation. Another remarkable result is the appearance of an infinite series of conservation laws for the KdV equation with the corresponding set of conserved quantities [30, 31, 32], the so-called KdV integrals. We can apply these techniques to our master equations in the frequency domain and generate the associated set of KdV integrals [15]. One can see that the KdV equation constitutes an isospectral deformation of the master equations, in such a way that the transmission coefficients and the QNMs are preserved by the KdV deformation. Crucially, it was shown in [15] that the KdV integrals are invariant under Darboux transformations. That is, they are the same for all the potentials in the set of master equations.
@@ I.3 Hyperboloidal approach to BHPT
The use of hyperboloidal foliations has proved very successful in its application to both fundamental and applied problems in spacetime dynamics and BH scattering. It provides one of the main ingredients in key results such as the semi-global stability of Minkowski by Friedrich [36], BH stability results in cosmological scenarios [37, 38, 39], the characterisation of QNMs as eigenvalues of a non-selfadjoint operator [40, 41], the numerical stability of the exploration of solution resolution in [4], or the analysis of the spectral instability of BH QNMs [42, 43, 44, 45]. In recent years hyperbolodial foliations have become one of the main tools for the theoretical and numerical analysis of compact object sources of gravitational waves (cf. excellent discussions in [46, 47, 48, 49, 50, 51]).
A most important aspect of the hyperboloidal scheme in our present setting, namely focusing on the identification of the relevant algebraic structures in BH scattering, is that it provides a neat separation of bulk and asymptotic degrees of freedom. This feature manifests explicitly in the structure of the infinitesimal time generator of perturbations and it is the consequence of two ingredients: i) the choice of a “hyperboloidal time” $\tau$ related to the “Cauchy time” $t$ through an appropriate height function (i.e. $\tau=t+h(x)$) introducing the hyperboloidal foliation $\{\Sigma_{\tau}\}$ to cast the dynamics, and ii) a compactification of the hyperboloidal hypersurfaces $\Sigma_{\tau}$. The first element adapts the treatment to the relevant outgoing asymptotics, whereas the second one recasts such outgoing boundary conditions in terms of the regularity of the scattered field at the boundaries (the latter is akin  [^2] We thank J.A. Valiente-Kroon for insisting on this point. We refer to [52] for a discussion on the adaptation of Melrose’s “geometric scattering” to the hyperboloidal approach to spacetime dynamics. to the underlying philosophy in Melrose’s ‘geometric scattering’ [53]). Enforcing points i) and ii) casts, upon a first-order reduction in time, the wave dynamics of the master function $\phi$ in Schrödinger form
@@
The structure of $\mathbb{L}$ reflects the above-commented separation into bulk and asymptotic structures, respectively encoded in the ${\cal L}_{1}$ and ${\cal L}_{2}$ parts. On the one hand, the operator ${\cal L}_{1}$ stands as the structural counterpart in the hyperboloidal scheme of the Schrödinger operator in the Cauchy slicing setting of section I.2. Specifically, ${\cal L}_{1}$ is a (singular) Sturm-Liouville operator, namely a second-order elliptic operator that includes the effective scattering potential and provides the elements for the discussion of the background structures in linear scattering (namely, Darboux covariance and KdV structures). Its “singular” character, namely the vanishing at the boundaries of its principal as a consequence of the (‘Melrose’s) compactification of $\Sigma_{\tau}$, results in the absence of explicit boundary conditions that are now encoded in the structure of the operator $\mathbb{L}$ itself. On the other hand, the operator ${\cal L}_{2}$ is the structural piece encoding the outgoing nature of the boundary conditions of the scattering problem. That is, if ${\cal L}_{1}$ incorporates the boundary conditions into $\mathbb{L}$, it is ${\cal L}_{2}$ the element that characterises precisely its outgoing nature. It can actually be shown [42] that ${\cal L}_{2}$ is the term responsible of the non-self-adjoint character of $\mathbb{L}$ and that the corresponding dissipative character of the latter is precisely due to the flow of the scattered field through the spacetime boundaries, whose flux is controlled by the parameters in ${\cal L}_{2}$. Moreover, the term breaking the self-adjoint nature of the infinitesimal time generator, i.e. the operator $\mathbb{L}-\mathbb{L}^{\dagger}$ (where the adjoint $\mathbb{L}^{\dagger}$ is defined with respect to a natural ‘energy scalar product’) has a distributional nature with support at null infinity and the BH horizon. In other words, the missing DoFs in the hyperboloidal non-self-adjoint formulation of the problem, associated with the loss of DoFs propagating from the bulk through the boundary, are sharply accounted by the flux controlled by the ${\cal L}_{2}$ term, the latter being indeed a structure essentially living on the boundaries. This has led to the proposal in [54] that one should be able to account for the loss of “fast” DoFs reaching the boundaries in terms of formal DoFs “living” on the spacetime boundaries, that would work as a kind of “geiger counters” of “fast” DoFs (their inclusion in the problem rendering it conservative, since the total account of energy is conserved by construction). More ambitiously, Ref. [54] proposes that asymptotic symmetries [namely, the Bondi-Metzner-Sachs (BMS) group of symmetries [55, 56], or an appropriate extension] would stand as a dynamical symmetry generating the phase space of DoF on the boundary, namely the (asymptotic) “fast” DoFs.
@@ I.4 The “KdV-Virasoro-Schwarzian derivative” triangle: bulk structures in a BH dynamics program
Dwelling then in the bulk part, this work focuses on the bulk symmetries and explores the heuristic intuition that the KdV structure of BHPT [15, 16] is related to conformal symmetries of the bulk DoFs. This leads to the identification of the role of the Virasoro algebra in this setting and that of the Schwarzian derivative in the transformation of the scattering potential under conformal mappings. In this setting, the “KdV-Virasoro-Schwarzian derivative” triangle is identified as providing the basic ingredients for the underlying bulk integrability (cf. [57] for a systematic discussion in a larger integrability setting). This is done first in the setting of a Cauchy description, where the KdV symmetries first appeared [15, 16], then finding that such a “bulk” symmetry structure is (partially) preserved when passing to the hyperboloidal foliation.
The plan of the article is the following. In Sec. II we first review the basic concepts and tools of BHPT and then introduce the complete landscape of the possible master equations decouplings allowed by Einstein equations. In particular, in subsection II.3 we review the symmetry structure hiding behind such infinite hierarchy of master equations, namely Darboux covariance. In Sec. III we explore the bulk symmetries, in Cauchy slices, associated to KdV integrable structures. First, we introduce the KdV isospectrality and some interesting relations to SUSY quantum mechanics. Then, we show the connection with the conformal transformations and the Virasoro algebra. We conclude this section showing how the conformal transformations act on the master equations. In Sec. IV we present how the structures studied in the previous section can be carried to a hyperboloidal scheme. With this aim we review the hyperboloidal approach in BHPT and introduce a general transformation of the master function, showing the covariance of the hyperboloidal master equation under such transformation. Finally, we conclude with a specific choice of the transformation which provides the connection with the bulk symmetries investigated in Sec. III.
@@ II Linear scattering on a stationary background: BHPT
where one assumes that $|h_{\mu\nu}|\ll|\widehat{g}_{\mu\nu}|$ (see, e.g. [61, 62] for a precise formulation of this statement). Here, the background metric $\widehat{g}_{\mu\nu}$ satisfies the full Einstein equations while the dynamics of the perturbations $h_{\mu\nu}$ is linear. Schematically, this can be written as follows
@@
where $\delta G$ is the linearized Einstein operator. However, the identification of background/“slow” and perturbative/“fast” DoFs is not apparent at this stage, as we still have to identify the correct background quantities which are related to integrable/stable structures and the perturbative DoFs associated to the linear asymptotic dynamics. The perturbation theory framework guides us towards the decoupling of Einstein equations for the perturbations (7) into (infinite) master equations with a potential and a wave function which will be directly related to “slow” and “fast” DoFs respectively. This already reduces the problem to the study of linear scattering (of gravitational perturbations) on a stationary background  [^3] We however only consider static spherically symmetric backgrounds in this work..
@@ II.1 The frozen slow DoFs: spherically symmetric background
where $f(r)$ is given in Eq. (9) and we have chosen $x^{a}=(t,r)$ and $\Theta^{A}=(\theta,\varphi)$. In Eq. (10), $r(x^{a})$ is the areal radial coordinate and $r^{2}$ is the geometry warp factor. The covariant derivative on $S^{2}$ is denoted by a vertical bar ($\Omega_{AB|C}=0$) while for $M^{2}$ we use a colon ($g_{ab:c}=0$). The volume Levi-Civita tensor in $S^{2}$ is denoted by $\epsilon_{AB}$ while in $M^{2}$ is denoted by $\varepsilon_{ab}\,$. Something interesting about the metric of $M^{2}$ [see Eq. (11)] is that, if we use the tortoise coordinate
@@
It is important to stress that while the perturbative approach linearizes the dynamics of fast degrees of freedom, it provides a simplified but non-trivial framework to investigate the presence of (universal) integrable structure governing the slow/background DoFs. Indeed, while the dynamics
of perturbations is linear by construction [see Eq. (7)], the background is a fully nonlinear solution with a rich structure acting on the perturbations.
@@ II.2 Linear dynamics of the fast DoFs: master equations
We are not going to show explicit expressions for the harmonic components of the metric perturbations since they are not relevant in this work. They can be found, for instance, in Refs. [15, 59]. We also drop the harmonic indices from now on for the sake of clarity. The separation in Eq. (16) allows to decouple Einstein perturbative equations into an infinite set of equations for the different $(\ell,m)$ and parity modes (see, e.g. [65, 66]). Now, let us consider the Einstein perturbative equations for just a single mode, that is, for fixed harmonic numbers $(\ell,m)$ and parity. Moreover, we focus on the radiative modes, i.e. $\ell\geq 2$, neglecting, for the time being, the non-radiating modes $\ell=0,1$ [20] (see also [59]). These equations, seven for the even-parity case and three for the odd-parity case, are $1+1$ linear PDEs (i.e. only with dependence on the coordinates of $M^{2}$) that can be combined in such a way that we can isolate (decouple) special gauge-invariant combinations of the metric perturbations (including their first-order derivatives), known as master functions. In this way, the perturbative Einstein equations for a single mode reduce to wave-type equations with a potential, known as master equations. The concrete form of the master equations is
@@
where $x$ is the tortoise coordinate introduced in Eq. (13). In the wave equation (17), $\Psi^{\rm even/odd}(t,r)$ is the even/odd-parity master function; and $V^{\rm even/odd}_{\ell}(r)$ is the associated potential, which only depends on the harmonic number $\ell$ and the parity. For the cases where we have explicit expressions of the potentials, these are in terms of the areal radial coordinate $r$, but for the purposes of solving the master wave-type equations, we have to understand this dependence on $r$ as on a function of the radial coordinate, i.e. on $r=r(x)$.
@@
where $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are arbitrary constants. Therefore, for odd-parity perturbations, the master function is a linear combination of the Regge-Wheeler (RW) [19] and the Cunningham-Price-Moncrief (CPM) [67, 68, 69] master functions. In the case of even-parity perturbations, the most general master function is instead a linear combination of the Zerilli-Moncrief (ZM) [20, 70] and another one (NE) which was first found in covariant form in Ref. [14]  [^5] This master function already appeared in Ref. [86], in Schwarzschild coordinates and in the Regge-Wheeler gauge, as a better choice for computations in the context of the close limit approximation.. We are not going to show here the explicit covariant expressions as they can be found in Refs. [14, 18].